# Generating quantum non-local entanglement with mechanical rotations

### v1

@@ III Violation of the Bell-chsh inequality
Local realism enforces the Bell-CHSH inequality $|S|\leq 2$. However, for suitable choices of $\phi$, Eq. (11) allows to violate such constraint. In particular, as stated in Eq. (8), by setting $\phi=\pi/2+k\pi\leavevmode\nobreak\ (k\in\mathcal{Z})$, the state in Eq. (7) reduces to the Bell state $|\Psi^{-}\rangle$ and we achieve the notorious Tsirelson’s bound $|S|=2\sqrt{2}$ [41]. This occurs when the mechanical frequency $\Omega$ takes the values
Here, $k<0$ ($k>0$) would correspond to an (anti-) clockwise sense of rotation. Needless to say, the postselection process required to get Eq. (7) entails that only coincidence events should be considered in order to construct $S$. These occur with a overall detection probability
$P=\frac{1+\text{cos}(\phi)^{2}}{32}.$ (13)
In Fig. 1(b), we plot the Bell-CHSH function and such probability of detection for a set of values of the relevant physical parameters that are well within reach of existing photonic technology. We remark that photon losses would only lower the probability of detection, without affecting the quality of the resulting state achieved through our scheme, which is thus robust against the most relevant source of imperfections in our chosen experimental platform.
@@ IV Discussion
We have proposed a method for the controlled generation of non-locality using mechanical rotation that achieves the map
$|H\,V\rangle\rightarrow(\text{cos}(\phi)+1)|H\,V\rangle+(\text{cos}(\phi)-1)|V\,H\rangle,$ (14)
@@
where we have omitted the normalization for brevity (see Eqs. (1) and (7)), and we recall that $\phi\propto\Omega$ is the Sagnac phase proportional to the angular frequency of the mechanical rotation $\Omega$. By controlling the value of $\Omega$ we can thus prepare separable or nonlocally entangled final states.
@@
The map in Eq. (14) satisfies a number of desiderata: (i) For $\phi=0$ the transformation reduces to the identity map. In other words, without mechanical rotation, the state remains invariant (and classical). (ii) For $\phi=\pi\,k$ ($k\in{\mathbb{Z}})$ the transformation is either the identity operation (even $k$) or induces a polarization flip (odd $k$). The latter case shows that mechanical rotation can be used to swap the polarization state of photon pairs. (iii) For $\phi=(2k+1)\pi/2$ ($k\in{\mathbb{Z}})$ we generate the Bell state $|\Psi^{-}\rangle$. The generated state is expected to induce a maximal violation of the CHSH inequality given by the Tsirelson’s bound [41]. Importantly, the state $|\Psi^{-}\rangle$ arises only when we tune the mechanical frequency of rotation to the value $\Omega_{\text{Bell}}$ given in Eq. (12).
@@
The scheme is also robust against imperfections and noise due to the inherent protection characteristic of the Sagnac loop. Suppose some unwanted phases would be accumulating depending on the polarization $H$, $V$; this would contribute only to a global phase in Eq. (3), but no measurable differential phase would be generated. Similarly, any other random phase affecting the co-rotating path will automatically affect also the counter-rotating path, thus factoring out without affecting the final state. As shown in Fig. 1(b) the experimental parameters required to test the maximum violation of the CHSH inequality can be achieved with current photonic technologies by adaptation of previous experimental schemes [23, 28, 25, 26, 29, 24]. As such we do not expect any new fundamental or technical issue in the implementation of this proposal.
@@
A further benefit of the proposed scheme is also that it does not rely on specific models, but rather on the well established CHSH test. Naively, one would think that the post-selection step from Eq. (6) to Eq. (7) might be responsible for generating non-locality. However, this is not the case as without rotation, Eq. (7) reduces to a separable state. Hence non-zero mechanical rotation is a critical factor for generating non-locality in this setup, i.e., we can legitimately speak of rotationally induced nonlocality. Furthermore, it has been shown that in case of high-efficiency detectors, the violation of the CHSH inequality provides a measurement of nonlocality also in the frame of post-selection [43]. Other types of inequalities can be considered in the experiments to account for inefficient quantum detectors [43].
@@ Acknowledgements
MT acknowledges funding by the Leverhulme Trust (RPG-2020-197). MP acknowledges the support by the Horizon Europe EIC Pathfinder project QuCoM (Grant Agreement No. 101046973), the Leverhulme Trust Research Project Grant UltraQuTe (grant RGP-2018-266), the Royal Society Wolfson Fellowship (RSWF/R3/183013), the UK EPSRC (EP/T028424/1), and the Department for the Economy Northern Ireland under the US-Ireland R&D Partnership Programme. DF acknowledges support from the Royal Academy of Engineering and the UK EPSRC (EP/W007444/1).

### v2

@@ III.2 Tsirelson’s bound with postselection
In particular, as stated in Eq. (15), by setting $\phi=\pi/2+k\pi\leavevmode\nobreak\ (k\in\mathcal{Z})$, the state in Eq. (14) reduces to the Bell state $|\Psi^{-}\rangle$ and we achieve the notorious Tsirelson’s bound $|S|=2\sqrt{2}$ [41]. In Fig. 1(b), we plot the Bell-CHSH function defined in Eq. (16) for the same set of values as for the case without postselection to ease the comparison.
Saturating Tsirelson’s bound depends critically on two steps: (i) on the post-selection step from Eq. (6) to (14), where we have discarded the unfavorable state, and (ii) on a non-zero rotationally induced phase $\phi\neq 0$, which arises only for non-zero frequencies of rotation $\Omega\neq 0$. The post-selection step is however not enough to induce entanglement in the absence of mechanical rotation as discussed above for the case $\phi=0$. Furthermore, even without any form post-selection we have shown that the Bell-CHSH inequality remains violated (see Sec. III.1).
@@ IV Discussion
We have proposed a method for the controlled generation of quantum non-locality using mechanical rotation that achieves the map
$\hat{a}_{H}^{\dagger}\hat{b}_{V}^{\dagger}|0\rangle\rightarrow\frac{1}{2}\bigg{[}$ $(\text{cos}\phi-1)\hat{a}_{H}^{\dagger}\hat{b}_{V}^{\dagger}+(\text{cos}\phi+1)\hat{b}_{H}^{\dagger}\hat{a}_{V}^{\dagger}$
$+\text{sin$\phi$}\,(\hat{a}_{H}^{\dagger}\hat{a}_{V}^{\dagger}-\hat{b}_{H}^{\dagger}\hat{b}_{V}^{\dagger})\bigg{]}|0\rangle.$ (17)
@@
The minimalist derivation is fully contained in Eqs. (1)-(8), assuming only the Sagnac phase and the form of the beam-splitter transformation. By controlling the angular frequency of the mechanical rotation, $\Omega$, and hence the corresponding Sagnac phase, $\phi\equiv\phi(\Omega)$, we have shown that we can either prepare separable or nonlocally entangled final states.
@@
The map in Eq. (17) satisfies a number of desiderata: (i) For $\phi=0$ the transformation reduces to the identity map, i.e., $\mathcal{M}=\mathbb{I}$. In other words, without mechanical rotation, the state remains invariant (and classical). (ii) For $\phi=\pi\,k$ ($k\in{\mathbb{Z}})$ the transformation in Eq. (17) is either the identity operation (even $k$) or induces a polarization flip (odd $k$). The latter case shows that mechanical rotation can be used to swap the polarization state of photon pairs. (iii) For $\phi=(2k+1)\pi/2$ ($k\in{\mathbb{Z}})$ we obtain the state in Eq. (12), which is predicted to violate the Bell-CHSH inequality. The first line of Eq. (17) (corresponding to Alice and Bob detecting each one photon) reduces to the Bell state $|\Psi^{-}\rangle$, which is expected to induce a maximal violation of the Bell-CHSH inequality given by the Tsirelson’s bound [41]. However, even in absence of any form of post-selection, the total final state in Eq. (17) produces a violatation of the Bell-CHSH inequality up to the value $S=1+\sqrt{2}$ [42].
@@
The scheme is also robust against imperfections and noise due to the inherent protection characteristic of the Sagnac loop. Suppose some unwanted phases would be accumulating depending on the polarization $H$, $V$; this would contribute only to a global phase in Eq. (3), but no measurable differential phase would be generated. Similarly, any other random phase affecting the co-rotating path will automatically affect also the counter-rotating path, thus factoring out without affecting the final state. As shown in Fig. 1(b) the experimental parameters required to test the maximum violation of the Bell-CHSH inequality can be achieved with current photonic technologies by adaptation of previous experimental schemes [23, 27, 28, 25, 26, 29, 24]. As such we do not expect any new fundamental or technical issue in the implementation of this proposal (see Appendix C for the analysis of background noise, dark counts, and detector inefficiencies).
@@
A further benefit of the proposed scheme is also that it does not rely on specific models, but rather on the well established Bell-CHSH test. The violation arises only when we tune the mechanical frequency of rotation to the interval centered on the value $\Omega_{\text{Bell}}$. Hence non-zero mechanical rotation is a critical factor for generating non-locality in this setup, i.e., we can legitimately speak of rotationally induced nonlocality.
@@ Acknowledgements
MT acknowledges funding from the Slovenian Research and Innovation Agency (ARIS) under contracts N1-0392, P1-0416, SN-ZRD/22-27/0510 (RSUL Toroš) and from the Leverhulme Trust (RPG-2020-197). MP acknowledges the support by the Horizon Europe EIC Pathfinder project QuCoM (Grant Agreement No. 101046973), the Leverhulme Trust Research Project Grant UltraQuTe (grant RGP-2018-266), the Royal Society Wolfson Fellowship (RSWF/R3/183013), the UK EPSRC (EP/T028424/1), and the Department for the Economy Northern Ireland under the US-Ireland R&D Partnership Programme. DF acknowledges support from the Royal Academy of Engineering and the UK EPSRC (EP/W007444/1).
Appendix A Angular velocity fluctuations
The presented scheme is robust with respect to angular velocity fluctuations, $\delta\Omega$, of the rotating platform. For example, let us focus on the region of the first peak shown in Fig. 1, where $\phi\in[0,\pi]$. We can then readily estimate the minimum (maximum) angular frequency $\Omega_{-}$ ($\Omega_{+}$) at the boundaries of the interval, where we have a violation of the Bell inequality, i.e., determined by the condition $|S|>2$. We define the maximum allowed angular velocity fluctuations as $\delta\Omega\equiv\Omega_{+}-\Omega_{\text{Bell}}$ (or equivalently $\delta\Omega=\Omega_{\text{Bell}}-\Omega_{-}$ given the the symmetric shape of the peak), such that for angular frequencies in the interval
$\Omega\in\Omega_{\text{Bell}}\pm\delta\Omega$ (18)
we have a violation of the Bell inequality. From Eqs. (4) and  (11) (or from Eqs. (4) and (16)), we find the that the maximum allowed angular velocity fluctuations are given by
$\delta\Omega=\frac{c^{2}}{4A\omega}\arctan\sqrt{\frac{\sqrt{2}-1}{2}}.$ (19)
In particular, for the parameters listed in Fig. 1 we find the value $\delta\Omega=2\pi\times 0.1\text{Hz}$, which does not pose a challenging requirement for the rotation control of low-frequency rotating platforms [24, 28]. A similar analysis on the sensitivity of the Bell-CHSH violation can be performed also for the other peaks at higher ($k>0$) or negative ($k<0$) values of $\Omega_{\text{Bell}}$.
Appendix B Coincidence detection
The detection method considered here relies on four-photon coincidences: two photon pairs are generated by two different SPDC processes. One photon from each pair is heralded (not shown in Fig. 1(a)), ensuring that the other photon propagates within the system. In this frame, the loss of one photon do not affect the measurement; it will only result in longer acquisition times. In the following discussion, we do not explicitly include the heralding process; all references to two-photon detection or coincidence events are made under the implicit assumption that the corresponding heralded photons have been successfully detected.
A further point needs to be made: in the case with post-selection, not all two-photons detections produce an increase in the number of counts; only when the top-left pair of detectors (Alice) detects one photon and the bottom-right pair of detectors (Bob) detects one photon we are able to update the experimental value of $S$ defined in Eq. (16) (see Fig. 1(a)) - we need a coincidence measurement between Alice and Bob. We can find the resulting reduction in the number of counts by noting that the numerical prefactors of Eq. (6) and Eq. (7) give the associated probability amplitudes. Hence, squaring and summing the amplitudes of the terms $\propto\hat{a}^{\dagger}\hat{b}^{\dagger}$ we find the probability of detection $(\text{cos}^{2}(\phi)+1)/2$, which gives a $50\%$ reduction in the statistics at $\Omega=\Omega_{\text{Bell}}.$ The case without post-selection is unaffected by this reduction in the probability of detection as all the photons are taken into account.
Appendix C Effect of Noise on Bell-CHSH Test
Figure 2: Variation of the Bell-CHSH parameter S with frequency and error probability $\epsilon$ (a) Case with post-selection (b) Case without post-selection. We use same values for wavelength $\lambda$ and interferometric area as used in Fig. 1.
Noise in the form of background noise, dark counts, and detector inefficiencies can significantly affect the quantum correlations relevant to the Bell-CHSH test. In this section, we quantitatively examine the effect of such noise on the Bell-CHSH test and consequently on the extent of violation of local-realism. Such investigations are of paramount importance when trying to assess the feasibility of our scheme for experimental implementation.
One possible noise source is a passive coupling mechanism caused by back-scattering from the cavity mirrors, in particular the BS surfaces in our setup. While such coupling could, in principle, lead to spurious mode interactions or background contributions, it is not expected to play a significant role in our case. The single photons do not populate a well-defined intracavity mode, and therefore do not establish conditions for mode-locking through mirror-induced feedback. Furthermore, due to the low generation efficiencies, as well as the finite spectral bandwidth of the photons, their coherence length remains short. This prevents significant temporal or spatial overlap between co-rotating and counter-rotating photon wave packets, effectively suppressing coherent interference effects arising from back-scattering.
Low pair-generation efficiencies also play a crucial role in suppressing unwanted multi-pair emission events, which can introduce noise into the measurement outcomes. Similar to what occurs in static (non-rotating) measurements, such multi-pair events contribute to a small but non-negligible background. However, their impact can be minimized by operating at low pump powers, reducing the probability of higher-order emissions.
A separate treatment is necessary when considering a noise model in which the subsystems associated with Alice and Bob are subject to independent error processes with noise error probability $\epsilon_{A}$ and $\epsilon_{B}$, respectively [51]. To simplify the presentation and estimate the magnitude of the effects, we here assume that the noise error probability of Alice’s subsystem, $\epsilon_{A}$, is the same as Bob’s, $\epsilon_{B}$ ($\epsilon_{A}=\epsilon_{B}=\epsilon$), and that $0<\epsilon\ll 1$. Hence, we can write the total system statistical operator $\hat{\rho}$ as:
$\hat{\rho}=(1-2\epsilon)\hat{\rho}_{AB}+\epsilon\left(\hat{\rho}_{A}\otimes\frac{1}{2}\right)+\epsilon\left(\frac{1}{2}\otimes\hat{\rho}_{B}\right),$ (20)
where $\hat{\rho}_{AB}=\ket{\psi_{3}}\bra{\psi_{3}}$ |, with |$\ket{\psi_{3}}$ the interferometer output state give by Eq. (6), while $\hat{\rho}_{A}=\text{tr}_{B}\left(\ket{\psi_{3}}\bra{\psi_{3}}\right)$ and $\hat{\rho}_{B}=\text{tr}_{A}\left(\ket{\psi_{3}}\bra{\psi_{3}}\right)$ are the reduced states of Alice’s and Bob’s subsystem. We use $\hat{\rho}$ to compute the Bell-CHSH function $S$ from Eq. (9) for two cases: without post-selection (Sec. III.1) and with post-selection (Sec. III.2).
To evaluate E(a,b) in $S$, we use:
$E(\bm{a},\bm{b})=\text{tr}(\hat{O}\,\hat{\rho}),$ (21)
where $\hat{O}$ is equal to $(\bm{a}\cdot\bm{\sigma})\otimes(\bm{b}\cdot\bm{\sigma})$ for the case with post-selection (we have one photon for each subsystem). However, for the case without post-selection, we use the approach suggested in [42]: here $\hat{O}$ is equal to $(\bm{a}\cdot\bm{\sigma})\otimes(\bm{b}\cdot\bm{\sigma})$ when each subsystem receives one photon, while $\hat{O}$ is equal to identity for the other cases when Alice or Bob’s subsystem receive more than one photon each.
Using this formalism, for the case with post-selection, we get:
$S=(1-2\epsilon)4\sqrt{2}\frac{\sin^{2}\phi}{3+\cos{2\phi}}.$ (22)
Similarly for the case without post-selection, we get:
$S=(1-2\epsilon)(1+\sqrt{2)}\sin^{2}\phi+\epsilon\sin^{2}{\phi}.$ (23)
The variation of $S$ with frequency and the error probability $\epsilon$ is shown in Fig. 2. As expected, for the case with post-selection, we see a violation of the Bell-CHSH inequality for a higher error probability compared to the case without post-selection. Nevertheless, it is worth noticing that the Bell-CHSH parameter can exceed the threshold value $S=2$ in both cases. This is due to the fact that $S$ measures the nonlocality of the state and hence that our final state is indeed nonlocal, even if not maximally entangled.